# Stability for the 2-D plane Poiseuille flow in finite channel

### 31 Dec 2023 v1

@@ 2.2. The improved transition threshold with Navier-slip boundary condition
3. Space-time estimates for linearized Navier-Stokes equations around the Poiseuille flow: non-slip boundary condition
@@
$\left\{\begin{array}[]{lll}\partial_{t}\omega-\nu(\partial_{y}^{2}-k^{2})\omega+ik(1-y^{2})\omega+2ik\varphi=-ikf_{1}-\partial_{y}f_{2},\\ (\partial_{y}^{2}-k^{2})\varphi=\omega,\ \ u=(\partial_{y},-ik)\varphi,\\ \varphi(t,k,\pm 1)=\partial_{y}\varphi(t,k,\pm 1)=0,\\ \omega|_{t=0}=\omega_{0}(k,y),\end{array}\right.$ (3.1)
where $f_{i}\in L^{2}$ and $k\not=0$.
@@
Suppose the $\omega_{0}\in H^{1}_{0}$ with $\langle\omega_{0},e^{\pm|k|y}\rangle=0$, then the solution $\omega$ to (3.1) enjoys that the following estimates:
(1) If $\nu|k|^{2}\leq 2$, it holds that
$(\nu|k|)^{\frac{1}{2}}$ $\|\omega\|_{L^{2}L^{2}}^{2}+|k|(\|u\|_{L^{\infty}L^{2}}^{2}+\|u\|_{L^{2}L^{2}}^{2})+\nu^{\frac{3}{2}}\|\partial_{y}\omega\|_{L^{2}L^{2}}^{2}$ (3.2)
$+\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{\infty}L^{2}}^{2}+|k|\|(1-y^{2})\omega\|_{L^{2}L^{2}}^{2}$
$\lesssim$ $|k|^{-1}\|(\partial_{y},|k|)\omega_{0}\|_{L^{2}}^{2}+\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}.$
(2) If $\nu|k|^{2}\geq 2$, it holds that
$\|\omega\|_{L^{\infty}L^{2}}^{2}$ $+|k|^{2}(\|u\|_{L^{\infty}L^{2}}^{2}+\|u\|_{L^{2}L^{2}}^{2})+\nu\|(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}$ (3.3)
$\lesssim$ $\|\omega_{0}\|_{L^{2}}^{2}+\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}.$
@@
Indeed, it remains to give the space-time estimates of $\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{\infty}L^{2}}^{2}$ and $|k|\|(1-y^{2})\omega\|_{L^{2}L^{2}}^{2}$ for $\nu|k|^{2}\leq 2$, other estimates can be found in [9].
By testing (3.1) by $(1-y^{2})\omega$, we have
$\frac{1}{2}$ $\frac{\mathrm{d}}{\mathrm{d}t}\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\nu\|(1-y^{2})^{\frac{1}{2}}(\partial_{y},|k|)\omega\|_{L^{2}}^{2}+\nu\int_{-1}^{1}(-2y)\partial_{y}\omega\overline{\omega}\mathrm{d}y$
$+ik\|(1-y^{2})\omega\|_{L^{2}}^{2}=\langle-ikf_{1},\omega\rangle+\langle f_{2},\partial_{y}((1-y^{2})\omega)\rangle+2ik\langle\varphi,(1-y^{2})\omega\rangle,$
@@
$\frac{1}{2}$ $\frac{\mathrm{d}}{\mathrm{d}t}\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\nu\|(1-y^{2})^{\frac{1}{2}}(\partial_{y},|k|)\omega\|_{L^{2}}^{2}+|k|\|(1-y^{2})\omega\|_{L^{2}}^{2}$
$\leq$ $2\nu\|\omega\|_{L^{2}}\|\partial_{y}\omega\|_{L^{2}}+2|k|\|\varphi\|_{L^{2}}\|(1-y^{2})\omega\|_{L^{2}}$
$+|k|\|f_{1}\|_{L^{2}}\|(1-y^{2})\omega\|_{L^{2}}+\|f_{2}\|_{L^{2}}\|\partial_{y}((1-y^{2})\omega)\|_{L^{2}}$
$\leq$ $\nu^{\frac{1}{2}}\|\omega\|_{L^{2}}^{2}+\nu^{\frac{3}{2}}\|\partial_{y}\omega\|_{L^{2}}^{2}+\frac{1}{2}|k|\|(1-y^{2})\omega\|_{L^{2}}^{2}+4|k|\|\varphi\|_{L^{2}}^{2}$
$+\frac{1}{2}\nu|k|^{2}\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\frac{1}{2}(\nu|k|^{2})^{-1}\|(1-y^{2})^{\frac{1}{2}}f_{1}\|_{L^{2}}^{2}$
$+\frac{1}{2}\nu^{-1}\|f_{2}\|_{L^{2}}^{2}+\frac{1}{2}\nu\|(1-y^{2})^{\frac{1}{2}}\partial_{y}\omega\|_{L^{2}}^{2}$
or
$\frac{1}{2}$ $\frac{\mathrm{d}}{\mathrm{d}t}\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\nu\|(1-y^{2})^{\frac{1}{2}}(\partial_{y},|k|)\omega\|_{L^{2}}^{2}+|k|\|(1-y^{2})\omega\|_{L^{2}}^{2}$
$\leq$ $2\nu\|\omega\|_{L^{2}}\|\partial_{y}\omega\|_{L^{2}}+2|k|\|\varphi\|_{L^{2}}\|(1-y^{2})\omega\|_{L^{2}}$
$+|k|\|f_{1}\|_{L^{2}}\|(1-y^{2})\omega\|_{L^{2}}+\|f_{2}\|_{L^{2}}\|\partial_{y}((1-y^{2})\omega)\|_{L^{2}}$
$\leq$ $\nu^{\frac{1}{2}}\|\omega\|_{L^{2}}^{2}+\nu^{\frac{3}{2}}\|\partial_{y}\omega\|_{L^{2}}^{2}+\frac{1}{2}|k|\|(1-y^{2})\omega\|_{L^{2}}^{2}+4|k|\|\varphi\|_{L^{2}}^{2}$
$+\frac{1}{2}|k|\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\frac{1}{2}|k|\|(1-y^{2})^{\frac{1}{2}}f_{1}\|_{L^{2}}^{2}$
$+\frac{1}{2}\nu^{-1}\|f_{2}\|_{L^{2}}^{2}+\frac{1}{2}\nu\|(1-y^{2})^{\frac{1}{2}}\partial_{y}\omega\|_{L^{2}}^{2}.$
Therefore, we obtain that
$\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{\infty}L^{2}}^{2}+\nu\|(1-y^{2})^{\frac{1}{2}}(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}+|k|\|(1-y^{2})\omega\|_{L^{2}L^{2}}^{2}$
$\lesssim$ $\|\omega_{0}\|_{L^{2}}^{2}+\nu^{\frac{1}{2}}\|\omega\|_{L^{2}L^{2}}^{2}+\nu^{\frac{3}{2}}\|\partial_{y}\omega\|_{L^{2}L^{2}}^{2}+|k|^{-1}\|u\|_{L^{2}L^{2}}^{2}$
$+\min\{(\nu|k|^{2})^{-1},|k|\}\|f_{1}\|_{L^{2}L^{2}}^{2}+\nu^{-1}\|f_{2}\|_{L^{2}L^{2}}^{2}$
$\lesssim$ $|k|^{-1}\|(\partial_{y},|k|)\omega_{0}\|_{L^{2}}^{2}+\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}.$
@@
4. Space-time estimates for linearized Navier-Stokes equations around the Poiseuille flow: Navier-slip boundary condition
In this section, we will establish the space-time estimates for linearized Navier-Stokes equations as follows
$\left\{\begin{array}[]{lll}\partial_{t}\omega-\nu(\partial_{y}^{2}-k^{2})\omega+ik(1-y^{2})\omega+2ik\varphi=-ikf_{1}-\partial_{y}f_{2},\\ (\partial_{y}^{2}-k^{2})\varphi=\omega,\ \ u=(\partial_{y},-ik)\varphi,\\ \varphi(t,k,\pm 1)=\omega(t,k,\pm 1)=0,\\ \omega|_{t=0}=\omega_{0}(k,y),\end{array}\right.$ (4.1)
where $\omega_{0}(k,\pm 1)=0$ and $k\not=0$. It is noted that $\omega_{0}(k,\pm 1)=0$ ensures $\omega(t,k,\pm 1)=0$ for any $t\geq 0$.
Before giving the main result of this section, let us recall two lemmas from [9].
@@
For $f\in H^{1}_{0}(-1,1)$, it holds that
$\|f\|_{L^{2}}^{2}\leq\|(1-y^{2})^{\frac{1}{2}}f\|_{L^{2}}^{\frac{4}{3}}\|(1-y^{2})^{-1}f\|_{L^{2}}^{\frac{2}{3}}\leq C\|(1-y^{2})^{\frac{1}{2}}f\|_{L^{2}}^{\frac{4}{3}}\|\partial_{y}f\|_{L^{2}}^{\frac{2}{3}}.$
Lemma 4.2.
Suppose that $(\partial_{y}^{2}-|k|^{2})\varphi=w,\varphi(\pm 1)=0,|k|\neq 0$. Then it holds that
$2\left(\bigg{\|}\varphi^{\prime}+\dfrac{2y\varphi}{1-y^{2}}\bigg{\|}_{L^{2}}^{2}+|k|^{2}\|\varphi\|_{L^{2}}^{2}\right)\leq\langle(1-y^{2})w+2\varphi,w\rangle,$
$\|(\varphi^{\prime},|k|\varphi)\|_{L^{2}}^{2}+|k|^{2}\|\varphi\|_{L^{2}}^{2}\leq\big{(}8/|k|^{2}+1\big{)}\langle(1-y^{2})w+2\varphi,w\rangle.$
The space-time estimates for (4.1) are stated as follows.
Proposition 4.1.
Let $\omega$ be the solution to (4.1) with $\omega_{0}\in H^{1}_{0}(-1,1)$, and $f_{i}\in L^{2}(i=1,2)$, then there exists $0<c_{0}\ll 1$ such that

### 03 Mar 2024 v2

@@ 3. Resolvent estimates for Orr-Sommerfeld equation
4. Space-time estimates for linearized Navier-Stokes equations around the Poiseuille flow
@@
$\left\{\begin{array}[]{lll}\partial_{t}\omega-\nu(\partial_{y}^{2}-k^{2})\omega+ik(1-y^{2})\omega+2ik\varphi=-ikf_{1}-\partial_{y}f_{2}-f_{3}-f_{4},\\ (\partial_{y}^{2}-k^{2})\varphi=\omega,\ \ u=(\partial_{y},-ik)\varphi,\\ \varphi(t,k,\pm 1)=\omega(t,k,\pm 1)=0,\\ \omega|_{t=0}=\omega_{0}(k,y),\end{array}\right.$ (4.1)
where $f_{3}|_{y=\pm 1}=0$ and $k\not=0$.
@@
Suppose the $\omega_{0}\in H^{1}_{0}\cap H^{2}$, then the solution $\omega$ to (4.1) enjoys that the following estimates:
$\|\omega\|_{L^{\infty}L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}+(\nu|k|)^{\frac{1}{2}}\|\omega\|_{L^{2}L^{2}}^{2}+|k|\|u\|_{L^{2}L^{2}}^{2}$ (4.2)
$\lesssim$ $\|\Delta_{k}\omega_{0}\|_{L^{2}}^{2}+\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}$
$+|k|^{-1}\|(\partial_{y},|k|)f_{3}\|_{L^{2}L^{2}}^{2}+\min\{(\nu|k|)^{-\frac{1}{2}},(\nu|k|^{2})^{-1}\}\|f_{4}\|_{L^{2}L^{2}}^{2}.$
Moreover, there exists $0<c\ll 1$ such that
$\|e^{c\nu^{\frac{1}{2}}t}\omega\|_{L^{\infty}L^{2}}^{2}+\nu\|e^{c\nu^{\frac{1}{2}}t}(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}$ (4.3)
$+(\nu|k|)^{\frac{1}{2}}\|e^{c\nu^{\frac{1}{2}}t}\omega\|_{L^{2}L^{2}}^{2}+|k|\|e^{c\nu^{\frac{1}{2}}t}u\|_{L^{2}L^{2}}^{2}$
$\lesssim$ $\|\Delta_{k}\omega_{0}\|_{L^{2}}^{2}+\nu^{-1}\|e^{c\nu^{\frac{1}{2}}t}(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}+|k|^{-1}\|e^{c\nu^{\frac{1}{2}}t}(\partial_{y},|k|)f_{3}\|_{L^{2}L^{2}}^{2}$
$+\min\{(\nu|k|)^{-\frac{1}{2}},(\nu|k|^{2})^{-1}\}\|e^{c\nu^{\frac{1}{2}}t}f_{4}\|_{L^{2}L^{2}}^{2}.$
To establish Proposition 4.1, let us introduce
$\omega=\omega_{I}+\omega_{H}$
with
$\left\{\begin{array}[]{lll}\partial_{t}\omega_{I}-\nu(\partial_{y}^{2}-k^{2})\omega_{I}+ik(1-y^{2})\omega_{I}+2ik\varphi_{I}=-ikf_{1}-\partial_{y}f_{2}-f_{3}-f_{4},\\ (\partial_{y}^{2}-k^{2})\varphi_{I}=\omega_{I},\ \ u_{I}=(\partial_{y},-ik)\varphi_{I},\\ \varphi_{I}(t,k,\pm 1)=\omega_{I}(t,k,\pm 1)=0,\\ \omega_{I}|_{t=0}=0,\end{array}\right.$ (4.4)
and
$\left\{\begin{array}[]{lll}\partial_{t}\omega_{H}-\nu(\partial_{y}^{2}-k^{2})\omega_{H}+ik(1-y^{2})\omega_{H}+2ik\varphi_{H}=0,\\ (\partial_{y}^{2}-k^{2})\varphi_{H}=\omega_{H},\ \ u_{H}=(\partial_{y},-ik)\varphi_{H},\\ \varphi_{H}(t,k,\pm 1)=\omega_{H}(t,k,\pm 1)=0,\\ \omega_{H}|_{t=0}=\omega_{0}.\end{array}\right.$ (4.5)
The space-time estimates for $\omega_{I}$ are stated as follows.
Lemma 4.1.
Let $\omega_{I}$ be the solution to (4.4), then it holds that
$\|\omega_{I}\|_{L^{\infty}L^{2}}^{2}+(\nu|k|)^{\frac{1}{2}}\|\omega_{I}\|_{L^{2}L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega_{I}\|_{L^{2}L^{2}}^{2}+|k|\|u_{I}\|_{L^{2}L^{2}}^{2}$
$\lesssim$ $\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}+|k|^{-1}\|(\partial_{y},|k|)f_{3}\|_{L^{2}}^{2}+\min\{(\nu|k|)^{-\frac{1}{2}},(\nu|k|^{2})^{-1}\}\|f_{4}\|_{L^{2}L^{2}}^{2}.$
In addition, there exists $0<c\ll 1$ such that
$\|e^{c\nu^{\frac{1}{2}}t}\omega_{I}\|_{L^{\infty}L^{2}}^{2}+(\nu|k|)^{\frac{1}{2}}\|e^{c\nu^{\frac{1}{2}}t}\omega_{I}\|_{L^{2}L^{2}}^{2}$ (4.6)
$\qquad+\nu\|e^{c\nu^{\frac{1}{2}}t}(\partial_{y},|k|)\omega_{I}\|_{L^{2}L^{2}}^{2}+|k|\|e^{c\nu^{\frac{1}{2}}t}u_{I}\|_{L^{2}L^{2}}^{2}$
$\lesssim$ $\nu^{-1}\|e^{c\nu^{\frac{1}{2}}t}(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}$
$+|k|^{-1}\|e^{c\nu^{\frac{1}{2}}t}(\partial_{y},|k|)f_{3}\|_{L^{2}}^{2}+\min\{(\nu|k|)^{-\frac{1}{2}},(\nu|k|^{2})^{-1}\}\|e^{c\nu^{\frac{1}{2}}t}f_{4}\|_{L^{2}L^{2}}^{2}.$
@@
Note that the proof is based on the resolvent estimates for $\mathcal{L}_{k}-\epsilon(\nu|k|)^{\frac{1}{2}}$, we only give the proof for the estimates without $e^{c\nu^{\frac{1}{2}}t}$, and the case with the weight can be obtained via perturbation argument.
By Proposition 3.1, we have
$(\nu|k|)^{\frac{1}{2}}\|\omega_{I}\|_{L^{2}L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega_{I}\|_{L^{2}L^{2}}^{2}$
$\lesssim$ $\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}+|k|^{-1}\|(\partial_{y},|k|)f_{3}\|_{L^{2}}^{2}$
$+\min\{(\nu|k|)^{-\frac{1}{2}},(\nu|k|^{2})^{-1}\}\|f_{4}\|_{L^{2}L^{2}}^{2},$
where the following energy estimates have been used
$\nu\|(\partial_{y},|k|)\omega_{I}\|_{L^{2}L^{2}}^{2}$
$\lesssim$ $\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}+|k|^{-1}\|(\partial_{y},|k|)f_{3}\|_{L^{2}}^{2}+(\nu|k|^{2})^{-1}\|f_{4}\|_{L^{2}L^{2}}^{2}$
and
$\nu|k|^{2}\|\omega_{I}\|_{L^{2}L^{2}}^{2}$
$\lesssim$ $\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}+|k|^{-1}\|(\partial_{y},|k|)f_{3}\|_{L^{2}}^{2}+\|f_{4}\|_{L^{2}L^{2}}\|\omega\|_{L^{2}L^{2}}$
$\lesssim$ $\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}+|k|^{-1}\|(\partial_{y},|k|)f_{3}\|_{L^{2}}^{2}+(\nu|k|)^{-\frac{1}{2}}\|f_{4}\|_{L^{2}L^{2}}^{2}.$
It remains to bound $\|u_{I}\|_{L^{2}L^{2}}$ and $\|\omega_{I}\|_{L^{\infty}L^{2}}$. Let us introduce
$\omega_{I}=\omega_{I}^{(1)}+\omega_{I}^{(2)}$
with
$\left\{\begin{array}[]{lll}\partial_{t}\omega_{I}^{(1)}-\nu(\partial_{y}^{2}-k^{2})\omega_{I}^{(1)}+ik(1-y^{2})\omega_{I}^{(1)}+2ik\varphi_{I}^{(1)}=-ikf_{1}-\partial_{y}f_{2}-f_{3},\\ \partial_{t}\omega_{I}^{(2)}-\nu(\partial_{y}^{2}-k^{2})\omega_{I}^{(2)}+ik(1-y^{2})\omega_{I}^{(2)}+2ik\varphi_{I}^{(2)}=f_{4},\\ (\partial_{y}^{2}-k^{2})\varphi_{I}^{(i)}=\omega_{I}^{(i)},\ \ u_{I}^{(i)}=(\partial_{y},-ik)\varphi_{I}^{(i)},\\ \varphi_{I}^{(i)}(t,k,\pm 1)=\omega_{I}^{(i)}(t,k,\pm 1)=0,\ \omega_{I}^{(i)}|_{t=0}=0,\end{array}\right.$ (4.7)
By Proposition 3.1, we have
$|k|\|u_{I}^{(1)}\|_{L^{2}L^{2}}^{2}\lesssim$ $\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}+|k|^{-1}\|(\partial_{y},|k|)f_{3}\|_{L^{2}}^{2}$
and
$|k|\|u_{I}^{(2)}\|_{L^{2}L^{2}}^{2}+(\nu|k|)^{\frac{1}{2}}\|\omega_{I}^{(2)}\|_{L^{2}L^{2}}^{2}\lesssim(\nu|k|)^{-\frac{1}{2}}\|f_{4}\|_{L^{2}L^{2}}^{2}.$
Meanwhile, one has
$|k|\|u_{I}^{(2)}\|_{L^{2}L^{2}}^{2}\leq|k|^{-1}\|\omega_{I}^{(2)}\|_{L^{2}L^{2}}^{2}\lesssim(\nu|k|^{2})^{-1}\|f_{4}\|_{L^{2}L^{2}}^{2}.$
Thus, it holds that
$|k|\|u_{I}^{(2)}\|_{L^{2}L^{2}}^{2}\lesssim\min\{(\nu|k|)^{-\frac{1}{2}},(\nu|k|^{2})^{-1}\}\|f_{4}\|_{L^{2}L^{2}}^{2}.$
Putting the above estimates together, the estimate for $\|u_{I}\|_{L^{2}L^{2}}$ follows.
Finally, we give the estimate for $\|\omega_{I}\|_{L^{\infty}L^{2}}$. Indeed, we have
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\omega_{I}\|_{L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega_{I}\|_{L^{2}}^{2}=\mathrm{Re}\langle-ikf_{1}-\partial_{y}f_{2}-f_{3}-f_{4},\omega_{I}\rangle$
$\leq$ $\|(f_{1},f_{2})\|_{L^{2}}\|(\partial_{y},|k|)\omega_{I}\|_{L^{2}}+\|(\partial_{y},|k|)f_{3}\|_{L^{2}}\|(\partial_{y},|k|)\varphi_{I}\|_{L^{2}}+\|f_{4}\|_{L^{2}}\|\omega_{I}\|_{L^{2}}$
$\leq$ $\frac{1}{2}\|(\partial_{y},|k|)\omega_{I}\|_{L^{2}}^{2}+\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}}^{2}+|k|^{-1}\|(\partial_{y},|k|)f_{3}\|_{L^{2}}^{2}+|k|\|(\partial_{y},|k|)\varphi_{I}\|_{L^{2}}^{2}$
$+(\nu|k|^{2})^{-1}\|f_{4}\|_{L^{2}}^{2}$
or
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\omega_{I}\|_{L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega_{I}\|_{L^{2}}^{2}=\mathrm{Re}\langle-ikf_{1}-\partial_{y}f_{2}-f_{3}-f_{4},\omega_{I}\rangle$
$\leq$ $\|(f_{1},f_{2})\|_{L^{2}}\|(\partial_{y},|k|)\omega_{I}\|_{L^{2}}+\|(\partial_{y},|k|)f_{3}\|_{L^{2}}\|(\partial_{y},|k|)\varphi_{I}\|_{L^{2}}+\|f_{4}\|_{L^{2}}\|\omega_{I}\|_{L^{2}}$
$\leq$ $\frac{1}{2}\|(\partial_{y},|k|)\omega_{I}\|_{L^{2}}^{2}+\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}}^{2}+|k|^{-1}\|(\partial_{y},|k|)f_{3}\|_{L^{2}}^{2}+|k|\|(\partial_{y},|k|)\varphi_{I}\|_{L^{2}}^{2}$
$+\frac{1}{2}\left((\nu|k|)^{-1/2}\|f_{4}\|_{L^{2}}^{2}+(\nu|k|)^{\frac{1}{2}}\|\omega_{I}\|_{L^{2}}^{2}\right),$
@@
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\omega_{I}\|_{L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega_{I}\|_{L^{2}}^{2}$
$\lesssim$ $\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}}^{2}+|k|^{-1}\|(\partial_{y},|k|)f_{3}\|_{L^{2}}^{2}$
$+\min\{(\nu|k|)^{-1/2},(\nu|k|^{2})^{-1}\}\|f_{4}\|_{L^{2}}^{2}+|k|\|u_{I}\|_{L^{2}}^{2}+(\nu|k|)^{\frac{1}{2}}\|\omega_{I}\|_{L^{2}}^{2}.$
Therefore, we have
$\|\omega_{I}\|_{L^{\infty}L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega_{I}\|_{L^{2}L^{2}}^{2}$
$\lesssim$ $\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}+|k|^{-1}\|(\partial_{y},|k|)f_{3}\|_{L^{2}L^{2}}^{2}$
$+\min\{(\nu|k|)^{-1/2},(\nu|k|^{2})^{-1}\}\|f_{4}\|_{L^{2}L^{2}}^{2}+|k|\|u_{I}\|_{L^{2}L^{2}}^{2}+(\nu|k|)^{\frac{1}{2}}\|\omega_{I}\|_{L^{2}L^{2}}^{2}$
$\lesssim$ $\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}+|k|^{-1}\|(\partial_{y},|k|)f_{3}\|_{L^{2}L^{2}}^{2}$
$+\min\{(\nu|k|)^{-1/2},(\nu|k|^{2})^{-1}\}\|f_{4}\|_{L^{2}L^{2}}^{2}.$
@@
Next we give the estimates for $\omega_{H}$.
@@
It holds that
[…]
$+(\nu|k|)^{\frac{1}{2}}\|\omega_{H}\|_{L^{2}L^{2}}^{2}+|k|\|u_{H}\|_{L^{2}L^{2}}^{2}\lesssim\|\Delta_{k}\omega_{0}\|_{L^{2}}^{2}.$
In addition, there exists $0<c\ll 1$ such that
$\|e^{c\nu^{\frac{1}{2}}t}$ $\omega_{H}\|_{L^{\infty}L^{2}}^{2}+(\nu|k|)^{\frac{1}{2}}\|e^{c\nu^{\frac{1}{2}}t}\omega_{H}\|_{L^{2}L^{2}}^{2}$
$+\nu\|e^{c\nu^{\frac{1}{2}}t}(\partial_{y},|k|)\omega_{H}\|_{L^{2}L^{2}}^{2}+|k|\|e^{c\nu^{\frac{1}{2}}t}u_{H}\|_{L^{2}L^{2}}^{2}\lesssim\|\Delta_{k}\omega_{0}\|_{L^{2}}^{2}.$
Proof.
The proof is based on the resolvent estimates for $\mathcal{L}_{k}-\epsilon(\nu|k|)^{\frac{1}{2}}$, we only give the proof for the estimates without $e^{c\nu^{\frac{1}{2}}t}$, and the case with the weight can be obtained via perturbation argument with standard energy argument.
By the Gearhart-Prüss type lemma in [43] and the resolvent estimates in Proposition 3.1, we have
$\|\omega_{H}(t)\|_{L^{2}}\leq Ce^{-c_{1}(\nu|k|)^{\frac{1}{2}}t}\|\omega_{0}\|_{L^{2}},$
which gives that
$\|\omega_{H}\|_{L^{\infty}L^{2}}^{2}+(\nu|k|)^{\frac{1}{2}}\|\omega_{H}\|_{L^{2}L^{2}}^{2}\lesssim\|\omega_{0}\|_{L^{2}}^{2}.$
The energy argument gives that
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\omega_{H}\|_{L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega_{H}\|_{L^{2}}^{2}=0,$
which yields that
$\nu\|(\partial_{y},|k|)\omega_{H}\|_{L^{2}L^{2}}^{2}\leq\|\omega_{0}\|_{L^{2}}^{2}.$
Next we give the estimate for $u_{H}$.
If $\nu|k|^{2}\leq 2$, let $\omega_{H}^{(0)}$ be the solution to
$\partial_{t}\omega_{H}^{(0)}+ik(1-y^{2})\omega_{H}^{(0)}+2ik\varphi_{H}^{(0)}=0,$
$(\partial_{y}^{2}-|k|^{2})\varphi_{H}^{(0)}=\omega_{H}^{(0)},\varphi_{H}^{(0)}|_{y=\pm 1}=0,\ \omega_{H}^{(0)}|_{t=0}=\omega_{0},$
due to $\omega_{0}\in H^{1}_{0}$, it holds that $\omega_{H}^{(0)}|_{y=\pm 1}=0$ for any $t\geq 0$.
Following the arguments of [9], let us introduce
$\omega_{H}^{(1)}=g(t,y)\omega^{(0)}_{H},\ g(t,y)=e^{-\nu|k|^{2}t-4\nu|k|^{2}y^{2}t^{3}/3-(\nu|k|)^{1/2}t},$
$(\partial_{y}^{2}-|k|^{2})\varphi_{H}^{(i)}=\omega_{H}^{(i)},\quad\varphi_{H}^{(i)}(\pm 1)=0,\quad i=0,1.$
It is easy to find that
$\partial_{t}\omega_{H}^{(1)}+\nu|k|^{2}\omega_{H}^{(1)}+(4\nu|k|^{2}y^{2}t^{2}+(\nu|k|)^{1/2})\omega_{H}^{(1)}$
$\qquad\qquad+ik(1-y^{2})\omega_{H}^{(1)}=-2ikg(t,y)\varphi_{H}^{(0)},$
$\omega_{H}^{(0)}|_{t=0}=\omega_{H}^{(1)}|_{t=0}=\omega_{0},\ \omega_{H}^{(1)}|_{y=\pm 1}=0.$
Thus, we have
$\partial_{t}\omega_{H}^{(1)}-\nu(\partial_{y}^{2}-|k|^{2})\omega_{H}^{(1)}+ik(1-y^{2})\omega_{H}^{(1)}+2ik\varphi_{H}^{(1)}$
$=-\nu\partial_{y}^{2}\omega_{H}^{(1)}-4\nu|k|^{2}y^{2}t^{2}\omega_{H}^{(1)}-(\nu|k|)^{1/2}\omega_{H}^{(1)}-2ik(g\varphi_{H}^{(0)}-\varphi_{H}^{(1)}).$
Now we decompose $\omega_{H}$ as follows
$\omega_{H}=\omega_{H}^{(1)}+\omega_{H}^{(2)},$
where $\omega_{H}^{(2)}$ satisfies